Commerce Corporation has a high probability of operating at 42,000 activity hours during the upcoming period, and lower probabilities of operating at 32,000 hours and 52,000 hours. The company's flexible budget revealed the following: 32,000 Hours 42,000 Hours 52,000 Hours Variable costs $ 176,000 $ 231,000 $ 286,000 Fixed costs 920,000 920,000 920,000 If Commerce operated at 37,000 hours, its total budgeted cost would be:
Answer #1

Answer: $ 1,123,500

Total Budgeted Cost at 37,000 Hours

Budgeted Cost Formula = Variable Cost per unit * Hours + Fixed cost

Varaible cost Per hour = 176,000/ 32,000 = $ 5.50 Per hour

Budgeted Cost = $5.50*37,000 Hours+ 920,000 = $ 1,123,500 (Answer)